OpenStudy (anonymous):

factoring out; -16t^2+64t-60=0

hartnn (hartnn):

factor out -4 first, what u get ?

OpenStudy (anonymous):

-4(4t^2-16t+15)=0

OpenStudy (skullpatrol):

(?t -?)(?t -?) fill in the ?'s

OpenStudy (anonymous):

(t-1)(t-15)=0

OpenStudy (anonymous):

t=1 t=15

hartnn (hartnn):

nopes, u need 2 numbers whose product is 15*4 = 60 and sum = -16 its easy, both numbers are negative

OpenStudy (anonymous):

hmm.. I am confused now not sure.

hartnn (hartnn):

-4(4t^2-16t+15)=0 -4(4t^2 - ? t-?t +15) = 0 what should replace those 2 ?'s two numbers whose sum is -16 and product = 60

OpenStudy (skullpatrol):

@malibugranprix2000 What numbers when multiplied together equal 60?

hartnn (hartnn):

if u think a bit, its really easy

OpenStudy (anonymous):

well there is 10 and 6

hartnn (hartnn):

thats correct -4(4t^2 - 10 t-6t +15) = 0 can u factor further ?

OpenStudy (anonymous):

not sure how

hartnn (hartnn):

ok, -4(4t^2 - 10 t-6t +15) = 0 -4(2t(2t-5)-3(2t-5)) = 0 u see what i did there ?

OpenStudy (anonymous):

oh I see

hartnn (hartnn):

what u think should be next step ?

OpenStudy (anonymous):

so is 2t-5=0, 2t-5=0 is taken out

OpenStudy (skullpatrol):

-4(4t^2-16t+15) = -4(2t-3)(2t-5) = 0

hartnn (hartnn):

yes, but we have other factor also, we write it like this -4(2t(2t-5)-3(2t-5)) = 0 -4(2t-5)(2t-3)=0

OpenStudy (skullpatrol):

2t-3=0 or 2t-5=0

hartnn (hartnn):

it was asked to factor.... 3 factors -4,(2t-5),(2t-3)

OpenStudy (skullpatrol):

{3/2,5/2}

OpenStudy (anonymous):

I see now thanks a lot.

hartnn (hartnn):

welcome ^_^

OpenStudy (skullpatrol):

Don't forget to check the solutions in the original equation. @malibugranprix2000

OpenStudy (anonymous):

thanks @hartnn and @skullpatrol